# Degenerate parabolic equations in divergence form: fundamental solution and Gaussian bounds

### v2

@@ Abstract.
In this paper, we consider second order degenerate parabolic equations with complex, measurable, and time-dependent coefficients. The degenerate ellipticity is dictated by a spatial $A_{2}$-weight. We prove that having a generalized fundamental solution with upper Gaussian bounds is equivalent to Moser’s $L^{2}$-$L^{\infty}$ estimates for local weak solutions. In the special case of real coefficients, Moser’s $L^{2}$-$L^{\infty}$ estimates are known, which provide an easier proof of Gaussian upper bounds, and a known Harnack inequality is then used to derive Gaussian lower bounds.
@@ 1. Introduction
The principal purposes of this paper is to establish the following theorem. Rigorous definitions are in Sections 2, 3 and 5.
@@
    if and only if all local weak solutions of $\mathcal{H}u=0$ and $\mathcal{H}^{\star}v=0$ satisfy Moser’s $L^{2}$-$L^{\infty}$ estimates. The function $\Gamma(t,x;s,y)$ is called a generalized fundamental solution.
@@
The existence and uniqueness of the fundamental solution family, proved in [AB24], are stated in Theorem 3.5, along with additional properties. The necessary and sufficient conditions of (1) are proven in Propositions 5.8 and 5.7, respectively. Finally, the special case of real-valued coefficients (2) is addressed in Section 6.
@@
Before delving into the details, it is useful to provide a brief survey of the results that have been obtained so far on this topic of fundamental solution of second order parabolic operators with Gaussian bounds, starting with the unweighted case, i.e. $\omega=1$. When the coefficients are regular, several methods exist for constructing the fundamental solution. The most effective technique combines a parametrix with the freezing point method [Fri08]. This method simplifies the problem by making the coefficients effectively independent of space, leading to explicit solutions represented by smooth kernels $\Gamma(t,x,s,y)$ with Gaussian decay. The treatment of the case of real, merely measurable coefficients was systematically addressed by Aronson [Aro67, Aro68]. He constructed generalized fundamental solutions and proved upper and lower bounds relying on the regularity properties of local weak solutions established by Nash [Nas58] and its extensions, alongside taking limits from operators with regular coefficients. In the case of complex coefficients, Auscher [Aus96] established Gaussian upper bounds estimates for the fundamental solutions of parabolic equations with complex coefficients, provided that these coefficients are time-independent and are small perturbations of real coefficients. Later, Hofmann and Kim [HK04] proved the equivalence between Moser’s $L^{2}$-$L^{\infty}$ estimates for solutions to parabolic systems and Gaussian upper bounds for the fundamental solution, extending Auscher’s result to include time-dependent coefficients and even systems. We note that it was assumed qualitatively that $A$ is smooth, which makes $\Gamma(t,x;s,y)$ available from literature.
@@
In the weighted case ($\omega\neq 1$), a work by Cruz-Uribe and Rios [CUR14] establishes the existence of the fundamental solution when $A$ is real-valued, symmetric and independent of $t$. It is given using the semi-group $(e^{-t\mathcal{L}})_{t>0}$, with $\mathcal{L}=-\omega^{-1}\mathrm{div}_{x}(A(x)\nabla_{x})$ and then exploit a Harnack inequality to derive Gaussian bounds. In a recent work, Ataei and Nyström [AN24] extended this result to real-valued and time-dependent coefficients using an approximation argument based on Kato’s work [Kat61], along with the Hölder continuity of weak solutions (i.e., Nash’s result for degenerate equations) to construct the fundamental solution and prove pointwise Gaussian upper bounds.
@@
In this paper, having at hand the existence of a fundamental solution family no matter what $A$ is, we can generalize [HK04] to weighted parabolic equations without assuming smoothness. That is, we prove that having upper bounds of the form (1.2) is equivalent to Moser’s $L^{2}$-$L^{\infty}$ for local weak solutions of $\mathcal{H}$ and $\mathcal{H}^{\star}$. When the coefficients are real, this provides a direct approach, even extending the result of [AN24], for Gaussian upper bounds, and the argument does not require the use of Nash’s regularity result. The latter is indeed valid, and we use it to derive Gaussian lower bounds. For further details, we refer the reader to the main text.
@@
Throughout the paper, we fix an integer $n\geq 1$. For any $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}$ and $r>0$, we set $Q_{r}(t,x):=(t-r^{2},t]\times B(x,r)$ and $Q_{r}^{\star}(t,x):=[t,t+r^{2})\times B(x,r)$ where $B(x,r)$ is the Euclidean ball of radius $r$ and center $x$. Thus, $Q_{r}(t,x)$ and $Q_{r}^{\star}(t,x)$ denote the usual forward and backward in time parabolic cylinders. For any $x\in\mathbb{R}^{n}$ and $r>0$, we set $\omega_{r}(x):=\omega(B(x,\sqrt{r})):=\int_{B(x,\sqrt{r})}\omega(y)\mathrm{d}y$.
@@
By convention, a notation $C(a,b,...)$ for a constant means that it depends only on $(a,b,...)$.
@@ 2.1. The weight and function spaces
(2.1) $[\omega]_{A_{2}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x\right)\left(\frac{1}{|Q|}\int_{Q}\omega^{-1}(x)\ \mathrm{d}x\right)<\infty,$
@@
During all this paper, $\omega=\omega(x)$ denotes a fixed weight belonging to the Muckenhoupt class $A_{2}(\mathbb{R}^{n},\mathrm{d}x)$. By definition (2.1), the weight $\omega^{-1}$ is also in the Muckenhoupt class $A_{2}(\mathbb{R}^{n},\mathrm{d}x)$ with $[\omega^{-1}]_{A_{2}}=[\omega]_{A_{2}}$. We introduce the measure $\mathrm{d}\omega:=\omega(x)\mathrm{d}x$ and if $E\subset\mathbb{R}^{n}$ a Lebesgue measurable set, we write $\omega(E)$ instead of $\int_{E}\mathrm{d}\omega$. We recall that $\omega_{r}(x)=\omega(B(x,\sqrt{r}))$ for any $x\in\mathbb{R}^{n}$ and $r>0$.
It follows from (2.1) that there exits constants $\eta\in(0,1)$ and $\beta>0$, depending only on $n$ and $[\omega]_{A_{2}}$, such that
@@
whenever $Q\subset\mathbb{R}^{n}$ is a cube and for all measurable sets $E\subset Q$. In particular, there exists a constant $D$, depending only on $n$ and $[\omega]_{A_{2}}$, called the doubling constant for $\omega$ such that
@@ 2.2. The distributional duality bracket
We define $-\Delta_{\omega}$ as the unbounded self-adjoint operator on $L^{2}_{\omega}(\mathbb{R}^{n})$ associated to the positive symmetric sesquilinear form on $H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})$ defined by
@@
Let $I\subset\mathbb{R}$ be an open set. For $f\in L^{1}_{\mathrm{loc}}(I;L^{2}_{\omega}(\mathbb{R}^{n})^{n})$, $g\in L^{1}_{\mathrm{loc}}(I;L^{2}_{\omega}(\mathbb{R}^{n}))$ and $\beta\in[0,1]$, we define $-\omega^{-1}\mathrm{div}_{x}(\omega f),(-\Delta_{\omega})^{\beta/2}g\in\mathcal{D^{\prime}}(I\times\mathbb{R}^{n})$ by setting for all $\varphi\in\mathcal{D}(I\times\mathbb{R}^{n})$,
@@ 3. Existence, uniqueness, representation and regularity results
We apply the method developed in [AB24] in this context by taking $T=\nabla_{x}:H^{1}_{\omega}(\mathbb{R}^{n})\to L^{2}_{\omega}(\mathbb{R}^{n})^{n}$ which is a closed and densely defined operator on $L^{2}_{\omega}(\mathbb{R}^{n})$. Moreover, it is injective because the measure $\mathrm{d}\omega$ has infinite mass by (2.2), and $\mathcal{D}(\mathbb{R}^{n})$ is dense in $D(T)=H^{1}_{\omega}(\mathbb{R}^{n})$ with respect to the graph norm. Furthermore, we have $T^{\star}T=-\Delta_{\omega}$, hence $S=(T^{\star}T)^{1/2}=(-\Delta_{\omega})^{1/2}$.
To connect the degenerate parabolic operator defined above with the one defined in [AB24, Section 6], we define, for all $t\in\mathbb{R}$, the sesquilinear form $B_{t}:H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})\rightarrow\mathbb{C}$ by setting
@@
In this section, we present the key results obtained using the method developed in [AB24] within the context of this paper.
@@
In [AB24], we have used sesquilinear brackets and we easily bring ourselves to the distributional duality bracket defined above as we have
@@ 3.1. A Lions’ type embedding with integral identities
We begin by presenting the following proposition, which improves the classical Lions embedding theorem [Lio57]. This result corresponds to [AB24, Theorem 2.1] in this concrete case.
@@ 3.2. Fundamental solution
We define the fundamental solution family as representing the inverse of the degenerate parabolic operator $\mathcal{H}=\partial_{t}-\omega^{-1}\mathrm{div}_{x}(A(t,\cdot)\nabla_{x})$ on ×$\mathbb{R}\times\mathbb{R}^{n}$.
@@
The operator $\mathcal{H}$ on ×$\mathbb{R}\times\mathbb{R}^{n}$ has a unique fundamental solution, as does $\mathcal{H}^{\star}$. Moreover, we have the following properties. Functions vanishing at infinity are denoted by $C_{0}$.
@@
The existence and uniqueness follow from Lemma 6.24 and Theorem 6.25 in [AB24]. As for the properties, we refer to Corollary 6.21 and Proposition 6.22 in the same paper. ∎
@@ 3.3. The Cauchy problem and the fundamental solution
Moreover, $u\in C([0,\mathfrak{T}];L^{2}_{\omega}(\mathbb{R}^{n}))$ with $u(0)=\psi$, $\lim_{t\to\infty}u(t)=0$ if $\mathfrak{T}=\infty$ (by convention, set $u(\infty)=0$), $t\mapsto\|u(t)\|^{2}_{2,\omega}$ is absolutely continuous on $[0,\mathfrak{T}]$ and we can write the energy equalities. Furthermore, we have $(-\Delta_{\omega})^{\alpha/2}u\in L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))$ for any $\alpha\in(0,1]$ with $r=\frac{2}{\alpha}\in[2,\infty)$ with
$\sup_{t\in[0,\mathfrak{T}]}\|u(t)\|_{2,\omega}+\|(-\Delta_{\omega})^{\alpha/2}u\|_{L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))}\leq C(\|f\|_{L^{2}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))}+\left\|g\right\|_{L^{\rho^{\prime}}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))}+\|\psi\|_{2,\omega}),$
@@
The case $\mathfrak{T}=\infty$ follows straightforwardly from combining Theorems 8.1 and 6.35 in [AB24]. For the case $\mathfrak{T}<\infty$, existence follows from restriction from the case $\mathfrak{T}=\infty$, and uniqueness is easily obtained using Proposition 3.3 along with (2.4). ∎
@@ 4.1. $L^{2}$-decay for the fundamental solution
For all $k\geq 0$, we have $\mathrm{dist}(\mathrm{supp}(f),C_{k}(R))\sim 2^{k}R$. Therefore, using Cauchy-Schwarz inequality, Proposition 4.1 and (2.2), we have for all $k\geq 0$,
@@
$\leq C\sqrt{\beta}\sqrt{\omega(B(0,R))}\times e^{-c\frac{4^{k}R^{2}}{t-s}}\times 2^{\frac{n}{4\eta}(k+1)},$
where $\beta,\eta>0$ are the structural constants appearing in (2.2). Therefore, $\int_{\mathbb{R}^{n}}|\Gamma(t,s)f|\ \mathrm{d}\omega<\infty.$

### v3

@@ Abstract.
In this paper, we consider second order degenerate parabolic equations with complex, measurable, and time-dependent coefficients. The degenerate ellipticity is dictated by a spatial $A_{2}$-weight. We prove that having a generalized fundamental solution with upper Gaussian bounds is equivalent to Moser’s $^{2}-LL^{\infty}$ estimates for local weak solutions. In the special case of real coefficients, Moser’s $^{2}-LL^{\infty}$ estimates are known, which provide an easier proof of Gaussian upper bounds, and a known Harnack inequality is then used to derive Gaussian lower bounds.
@@ 1. Introduction
The principal purpose of this paper is to establish the following theorem. Rigorous definitions are in Sections 2, 3 and 5.
@@
    if and only if all local weak solutions of $\mathcal{H}u=0$ and $\mathcal{H}^{\star}v=0$ satisfy Moser’s $^{2}-LL^{\infty}$ estimates. The function $\Gamma(t,x;s,y)$ is called a generalized fundamental solution.
@@
The existence and uniqueness of the fundamental solution family, proved in [AB25], are stated in Theorem 3.6, along with additional properties. The necessary and sufficient conditions of (1) are proven in Propositions 5.27 and 5.12, respectively. Finally, the special case of real-valued coefficients (2) is addressed in Section 6.
@@
Before delving into the details, it is useful to provide a brief survey of the results that have been obtained so far on this topic of fundamental solution of second order parabolic operators with Gaussian bounds, starting with the unweighted case, i.e. $\omega=1$. When the coefficients are regular, several methods exist for constructing the fundamental solution. The most effective technique combines a parametrix with the freezing point method [FRI64]. This method simplifies the problem by making the coefficients effectively independent of space, leading to explicit solutions represented by smooth kernels $\Gamma(t,x,s,y)$ with Gaussian decay. The treatment of the case of real, merely measurable coefficients was systematically addressed by Aronson [ARO67, ARO68]. He constructed generalized fundamental solutions and proved upper and lower bounds relying on the regularity properties of local weak solutions established by Nash [NAS58] and its extensions, alongside taking limits from operators with regular coefficients. In the case of complex coefficients, Auscher [AUS96] established Gaussian upper bounds estimates for the fundamental solutions of parabolic equations with complex coefficients, provided that these coefficients are time-independent and are small perturbations of real coefficients. Later, Hofmann and Kim [HK04] proved the equivalence between Moser’s $^{2}-LL^{\infty}$ estimates for solutions to parabolic systems and Gaussian upper bounds for the fundamental solution, extending Aronson result to include time-dependent coefficients and even systems.
@@
In the weighted case ($\omega\neq 1$), a work by Cruz-Uribe and Rios [CR14] establishes the existence of the fundamental solution when $A$ is real-valued, symmetric and independent of $t$. It is given using the semi-group $(e^{-t\mathcal{L}})_{t>0}$, with $\mathcal{L}=-\omega^{-1}\mathrm{div}_{x}(A(x)\nabla_{x})$ and then exploit a Harnack inequality to derive Gaussian bounds. In a recent work, Ataei and Nyström [AN25] extended this result to real-valued and time-dependent coefficients using an approximation argument based on Kato’s work [KAT61], along with the Hölder continuity of weak solutions (i.e., Nash’s result for degenerate equations) to construct the fundamental solution and prove pointwise Gaussian upper bounds.
@@
In this paper, having at hand the existence of a fundamental solution family no matter what $A$ is, we can generalize [HK04] to weighted parabolic equations without assuming smoothness. That is, we prove that having upper bounds of the form (1.3) is equivalent to Moser’s $^{2}-LL^{\infty}$ for local weak solutions of $\mathcal{H}$ and $\mathcal{H}^{\star}$. When the coefficients are real, this provides a direct approach, even extending the result of [AN25], for Gaussian upper bounds, and the argument does not require the use of Nash’s regularity result. The latter is indeed valid, and we use it to derive Gaussian lower bounds. For further details, we refer the reader to the main text.
@@
Throughout this paper, we adopt the following notation.
  1. $\blacklozenge$
    We fix an integer $n\geq 1$.
  2. $\blacklozenge$
    For any $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}$ and $r>0$, we set $Q_{r}(t,x):=(t-r^{2},t]\times B(x,r)$ and $Q_{r}^{\star}(t,x):=[t,t+r^{2})\times B(x,r)$ where $B(x,r)$ is the Euclidean ball of radius $r$ and center $x$. Thus, $Q_{r}(t,x)$ and $Q_{r}^{\star}(t,x)$ denote the usual forward and backward in time parabolic cylinders.
  3. $\blacklozenge$
    For any $x\in\mathbb{R}^{n}$ and $r>0$, we set $\omega_{r}(x):=\omega(B(x,\sqrt{r})):=\int_{B(x,\sqrt{r})}\omega(y)\mathrm{d}y$.
  4. $\blacklozenge$
@@ 5. $\blacklozenge$
  6. $\blacklozenge$
    By convention, the notation $C=C(a,b,\dots)$ for a constant means that $C\in(0,\infty)$ and depends only on $(a,b,\dots)$.
@@ 2.1. The weight and function spaces
(2.1) $\omega]_{A_{2}}:=\sup_{Q\subset\mathbb{R}^{n}}\left(\frac{1}{|Q|}\int_{Q}\omega(x)\ \mathrm{d}x\right)\left(\frac{1}{|Q|}\int_{Q}\omega^{-1}(x)\ \mathrm{d}x\right)<\infty,$
@@
During all this paper, $\omega=\omega(x)$ denotes a fixed weight belonging to the Muckenhoupt class $A_{2}(\mathbb{R}^{n},\mathrm{d}x)$. By definition (2.1), the weight $\omega^{-1}$ is also in the Muckenhoupt class $A_{2}(\mathbb{R}^{n},\mathrm{d}x)$ with $\omega^{-1}]_{A_{2}}=[\omega]_{A_{2}}$. We introduce the measure $\mathrm{d}\omega:=\omega(x)\mathrm{d}x$ and if $E\subset\mathbb{R}^{n}$ a Lebesgue measurable set, we write $\omega(E)$ instead of $\int_{E}\mathrm{d}\omega$. We recall that $\omega_{r}(x)=\omega(B(x,\sqrt{r}))$ for any $x\in\mathbb{R}^{n}$ and $r>0$.
It follows from (2.1) that there exits constants $\eta\in(0,1)$ and $\beta>0$, depending only on $n$ and $\omega]_{A_{2}}$, such that
@@
whenever $Q\subset\mathbb{R}^{n}$ is a cube and for all measurable sets $E\subset Q$. In particular, there exists a constant $D$, depending only on $n$ and $\omega]_{A_{2}}$, called the doubling constant for $\omega$ such that
@@ 2.2. The distributional duality bracket
We define $\Delta_{\omega}$ as the unbounded self-adjoint operator on $L^{2}_{\omega}(\mathbb{R}^{n})$ associated to the positive symmetric sesquilinear form on $H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})$ defined by
@@
Let $I\subset\mathbb{R}$ be an open set. For $f\in L^{1}_{\mathrm{loc}}(I;L^{2}_{\omega}(\mathbb{R}^{n})^{n})$, $g\in L^{1}_{\mathrm{loc}}(I;L^{2}_{\omega}(\mathbb{R}^{n}))$ and $\beta\in[0,1]$, we define $\omega^{-1}\mathrm{div}_{x}(\omega f),(-\Delta_{\omega})^{\beta/2}g\in\mathcal{D^{\prime}}(I\times\mathbb{R}^{n})$ by setting for all $\varphi\in\mathcal{D}(I\times\mathbb{R}^{n})$,
@@ 3. Existence, uniqueness, representation and regularity results
We apply the method developed in [AB25] in this context by taking $T=\nabla_{x}:H^{1}_{\omega}(\mathbb{R}^{n})\to L^{2}_{\omega}(\mathbb{R}^{n})^{n}$ which is a closed and densely defined operator on $L^{2}_{\omega}(\mathbb{R}^{n})$. Moreover, it is injective because the measure $\mathrm{d}\omega$ has infinite mass by (2.2), and $\mathcal{D}(\mathbb{R}^{n})$ is dense in $D(T)=H^{1}_{\omega}(\mathbb{R}^{n})$ with respect to the graph norm. Furthermore, we have $T^{\star}T=-\Delta_{\omega}$, hence $S=(T^{\star}T)^{1/2}=(-\Delta_{\omega})^{1/2}$.
To connect the degenerate parabolic operator defined above with the one defined in [AB25, Section 6], we define, for all $t\in\mathbb{R}$, the sesquilinear form $B_{t}:H^{1}_{\omega}(\mathbb{R}^{n})\times H^{1}_{\omega}(\mathbb{R}^{n})\rightarrow\mathbb{C}$ by setting
@@
In this section, we present the key results obtained using the method developed in [AB25] within the context of this paper.
@@
In [AB25], we have used sesquilinear brackets and we easily bring ourselves to the distributional duality bracket defined above as we have
@@ 3.1. A Lions’ type embedding with integral identities
We begin by presenting the following proposition, which improves the classical Lions embedding theorem [LIO57]. This result corresponds to [AB25, Theorem 2.1] in this concrete case.
@@ 3.2. Fundamental solution
We define the fundamental solution family as representing the inverse of the degenerate parabolic operator $\mathcal{H}=\partial_{t}-\omega^{-1}\mathrm{div}_{x}(A(t,\cdot)\nabla_{x})$ on $\mathbb{R}\times\mathbb{R}^{n}$.
@@
The operator $\mathcal{H}$ on $\mathbb{R}\times\mathbb{R}^{n}$ has a unique fundamental solution, as does $\mathcal{H}^{\star}$. Moreover, we have the following properties. Functions vanishing at infinity are denoted by $C_{0}$.
@@
The existence and uniqueness follow from Lemma 6.24 and Theorem 6.25 in [AB25]. As for the properties, we refer to Corollary 6.21 and Proposition 6.22 in the same paper. ∎
@@ 3.3. The Cauchy problem and the fundamental solution
Moreover, $u\in C([0,\mathfrak{T}];L^{2}_{\omega}(\mathbb{R}^{n}))$ with $u(0)=\psi$, $\lim_{t\to\infty}u(t)=0$ if $\mathfrak{T}=\infty$ (by convention, set $u(\infty)=0$), $t\mapsto\|u(t)\|^{2}_{2,\omega}$ is absolutely continuous on $0,\mathfrak{T}]$ and we can write the energy equalities. Furthermore, we have $(-\Delta_{\omega})^{\alpha/2}u\in L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))$ for any $\alpha\in(0,1]$ with $r=\frac{2}{\alpha}\in[2,\infty)$ with
$\sup_{t\in[0,\mathfrak{T}]}\|u(t)\|_{2,\omega}+\|(-\Delta_{\omega})^{\alpha/2}u\|_{L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))}\leq C(\|f\|_{L^{2}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n})^{n})}+\left\|g\right\|_{L^{\rho^{\prime}}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))}+\|\psi\|_{2,\omega}),$
@@
The case $\mathfrak{T}=\infty$ follows straightforwardly from combining Theorems 8.1 and 6.35 in [AB25]. For the case $\mathfrak{T}<\infty$, existence follows from restriction from the case $\mathfrak{T}=\infty$, and uniqueness is easily obtained using Proposition 3.4 along with (2.5). ∎
@@ 4.1. $L^{2}$-decay for the fundamental solution
For all $k\geq 0$, we have $\mathrm{dist}(\mathrm{supp}(f),C_{k}(R))\sim 2^{k}R$. Therefore, by applying the Cauchy–Schwarz inequality together with Proposition 4.1, we obtain, for all $k\geq 0$,
@@
$\leq C\sqrt{\omega(B(0,R))}\times e^{-c\frac{4^{k}R^{2}}{t-s}}\times D^{\frac{1}{2}(k+1)},$
where we have used the doubling property (2.3) in the last line. Therefore, it follows that
$\int_{\mathbb{R}^{n}}|\Gamma(t,s)f|\ \mathrm{d}\omega<\infty.$